# Counting Stacked Objects from Multi-View Images

### v1

@@ 3.4 Dataset generation
in 00000152_RGB0001, 00000152_RGB0010, 00000152_RGB0015
Figure 5: Multi-view images. We generate 30 views from arbitrary angles for each of the validation scenes in our synthetic dataset.
To train the volume usage estimation network $\Phi$ of Sec. 3.3, we require pairs of images and ground-truth volume usage percentage. To this end, we created a novel dataset of physically simulated stacked objects and we measure the exact fraction of volume being used by the objects. We used the ABC dataset [9], that comprises a wide variety of computer-assisted design (CAD) models, and we first discard models that are not watertight or that have multiple connected components. We normalized models to fit in a cube of side $0.05$. We then generated a virtual 3D scene with a container, and dropped an initial batch of 100 identical objects in that box. We repeated this step until the box is full, that is to say, the union of objects is still intersecting the space above the box even after the physical simulation has converged.
Once the physical simulation was complete, we numerically measured the percentage of volume used as well as the total number of objects in the container. Finally, we render multiple views of the container and objects from several angles to allow 3D reconstruction, as illustrated in Fig. 5. Crucially, this rendering is performed with a ray-tracing engine, and produces realistic images where objects at the bottom of the box appear darker due to ambient occlusion. For each object, a realistic appearance is randomly selected from a range of metallic and plastic materials.
We ran this physical simulation for over 2500 shapes from the ABC dataset, and isolate a subset of 100 shapes to use as a test set. This dataset allows us not only to train $\Phi$, but also to run our complete pipeline on the multi-view images in order to measure the accuracy of our count estimate $\mathcal{N}_{est}$, as performed in Sec. 4. It also includes ground-truth depth maps, which we employ in an ablation study in our experiments to assess the requirement of accurate depth maps of our method. The resulting dataset encompasses a wide variety of shapes and is illustrated in Fig. 3. We report additional statistics that highlight the diversity of the proposed dataset in Fig. 4.
4 Experiments
We evaluate our method in two ways: measuring the accuracy of 3D counting as a whole and of the volume estimation, in Sec. 4.3 and Sec. 4.2, respectively. These evaluations are performed over two datasets. The first comprises 100 scenes representing physically simulated shapes from the ABC dataset [9]. These scenes were isolated after their generation and were not seen during the training of our occupied volume network. The second is made of 20 real scenes that were captured with a regular smartphone’s RGB camera, and no additional sensor. These captures offer multiple views around stacks of objects in a container, lying flat on a table or still enclosed in their packaging. We measure the ground-truth number of units manually for all scenes below 1000, or infer it from the weight for even larger counts. Some of these captures can be seen in Fig. 1, and more examples are shown in our supplementary material.
NAE SRE MAE sMAPE
BMNet+ [22] 0.91 0.87 320.50 158.87
SAM[8]+CLIP[17] 0.73 0.61 259.22 102.77
Ours 0.22 0.09 79.48 27.65
Table 1: Counting evaluation on synthetic dataset. We evaluate on our proposed validation dataset of 100 scenes. Since the total count varies widely from 30 to around 1000 units across the simulated scenes, we report normalized metrics.
NAE SRE MAE sMAPE
BMNet+ [22] 0.98 1.00 1606.58 162.28
SAM[8]+CLIP[17] 0.95 0.99 1780.41 120.99
Ours 0.15 0.01 248.45 34.20
Table 2: Counting evaluation on real data. We capture 20 real scenes comprised of objects in stacks, packages, and various containers for evaluation on real-world applications.
MAE RMSE sMAPE $R^{2}$
@@ 4 Experiments
Table 3: Occupied volume estimation. We evaluate our method against two other depth-map based methods, and against the mean estimator which predicts the mean $\gamma$ of the training set (32.3%).
Figure 6: Partial view. Since our method is designed to be robust to restricted views of the objects to count, it is able to infer an accurate occupied volume ratio from small openings in packaging, such as in the rice scene (Est. 21649, G.T. 20063 (approx.), $\gamma_{est}=43.1\%$).
@@ 4.1 Metrics
We use several metrics to assess the accuracy of object counting and occupied volume estimation. The object counts vary significantly across scenes, ranging from 36 to 20063. Thus, the Mean Absolute Error (MAE)—defined as $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,$ where $y_{i}$ is the ground truth count and $\hat{y}_{i}$ is the predicted count for each scene—tends to amplify the importance of scenes with a high counts. To mitigate this, we also report normalized metrics. We use the Normalized Absolute Error (NAE) and Squared Relative Error (SRE), which scale errors relative to the ground truth. The NAE provides a measure of the absolute error normalized by the total ground truth count across scenes, while SRE emphasizes larger errors and penalizes significant deviations in high-count scenes. They are computed as
NAE $=\frac{\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|}{\sum_{i=1}^{n}y_{i}},\hskip 10.00002pt\text{SRE}$ $=\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}y_{i}^{2}}\;.$
We also report the Symmetric Mean Absolute Percentage Error (sMAPE), which offers a normalized percentage error. It is defined as
ensuring that errors are scaled symmetrically between the prediction and ground truth counts. Finally, the coefficient of determination, $R^{2}$, measures the proportion of variance in the ground-truth occupied volume ratio $\gamma$ explained by our predictions
4.2 3D counting
@@
To the best of our knowledge, there is no previous work on counting from multiple images which does not either implicitly assume all objects to be visible or require additional sensors such as LiDARs. Thus we compare our 3DC approach against BMNet+ [22]. It predicts a density map over all pixels of an image, and the estimated count is then inferred by summing over all pixels. Additionally, we compare against a combination of SAM [8] and CLIP [17], where SAM is used to generate a large number of masks from an input image and CLIP uses a set of negative and positive text prompts to identify masks that represent an object of interest. The final count is then taken to be the number of these masks.
Our method easily outperforms both baselines by a large margin, both on the synthetic (Tab. 3) and the real (Tab. 3) datasets. In Fig. 6, we show that our method can effectively count from a partial view of the objects to be counted. We further observe that the quality of our results varies little between the synthetic and real datasets, highlighting the generalization ability of our approach. Additionally, we display intermediate stages in Fig. 8, providing further insights into our results. For more visualizations over the complete dataset of real scenes, we refer the reader to our supplementary material.
Figure 7: Failure cases. For some highly complex shapes, our method fails to predict an accurate $\gamma_{est}$. They are represented in our real dataset, such as the scenes pasta (Est. 2025, G.T. 588, $\gamma_{est}=23.9\%$) and pins (Est. 100, G.T. 205, $\gamma_{est}=15.9\%$).
Figure 8: Intermediate results. We visualize the outputs of the different stages of our pipeline to provide further insights into their results.
4.3 Occupied Volume Estimation
We now study the performance of the occupied volume network by itself. There is little work on occupied volume estimation. For this reason, we implement additional baselines to evaluate the effectiveness of our approach, and evaluate them on our proposed synthetic dataset.
@@ 4.3 Occupied Volume Estimation
Since our hypothesis is that the depth map contains important information to predict this value, we define a first baseline that we will refer to as DepthExtrapolated. Given the top view of the container, we compute the maximal depth using a monocular depth estimator and use it to normalize the depth map. Then, we average the resulting values of the $K$ pixels, yielding the volume fraction estimate
$\gamma_{est}^{norm}=\frac{1}{K}\sum\frac{d_{i}}{d_{max}}$
@@
We observed that this first baseline was biased to predict values lower than expected. We thus defined a second one we dubbed DepthCorrected, which uses linear regression to correct $\gamma_{est}^{norm}$ into a new estimate $\gamma_{est}^{corrected}$. This provides a reasonable baseline that is able to model the observation that depth maps with high variance tend to correspond to low ratios of occupied volume. Finally, we also compare with the mean estimator, that predicts the mean percentage of $32.3\%$ occupied volume for all inputs.
@@
As can be seen in Tab. 3, our method outperforms these baselines by a significant margin. While DepthCorrected is more successful than other baselines, showing that the depth information is indeed meaningful for this task, it does not fully predict the occupied volume. We interpret this as a strong clue that the ratio $\frac{d_{i}}{d_{max}}$ alone is not enough to predict $\gamma$, and our proposed network successfully learns to extract meaningful information from the depth maps. We hypothesize that, thanks to the pre-trained foundation model, our volume network is able to capture additional geometric information such as the influence of concavities in the final occupation of volume.
@@ 4.4 Ablation study
NAE SRE MAE sMAPE
Ours ($\mathcal{T}-$, $\mathcal{V}-$) 0.22 0.09 79.48 27.65
Ours ($\mathcal{T}+$, $\mathcal{V}-$) 0.28 0.11 100.12 30.93
Ours ($\mathcal{T}+$, $\mathcal{V}+$) 0.31 0.12 111.04 35.92
Table 4: Ablation study on 3D counting. If ground-truth depth maps are used during training, it is indicated as $\mathcal{T}+$, and $\mathcal{T}-$ otherwise.Similarly, for evaluation purposes if ground-truth depth-maps are used during validation, we indicate it as $\mathcal{V}+$.
MAE RMSE sMAPE $R^{2}$
Ours ($\mathcal{T}-$, $\mathcal{V}-$) 0.06 0.07 29.18 0.79
Ours ($\mathcal{T}+$, $\mathcal{V}-$) 0.08 0.11 32.01 0.52
Ours ($\mathcal{T}+$, $\mathcal{V}+$) 0.10 0.13 37.35 0.32
Table 5: Ablation study on occupied volume estimation.
We also perform additional comparisons to evaluate the sensitivity of our approach to the depth maps produced by the monocular depth estimator. Since our proposed synthetic dataset has ground-truth depth maps both on the training images and the validation images, we perform several experiments.
Recall from Sec. 3.3 that, at training time, we use Depth Anything V2 [31] depth maps, a setting we refer to as $\mathcal{T}-$ in Tab. 3 and occupied volume estimation in Tab. 3. Instead, we could use ground-truth depth maps during training, a setting we will refer to as $\mathcal{T}+$, Similarly, for validation purposes, if ground-truth depth-maps are used we designate the method as $\mathcal{V}+$, and $\mathcal{V}-$ otherwise. Thus, the standard configuration of our method is $\mathcal{T}-,\mathcal{V}-$. The others are only used for ablation study purposes.
We report the results in Tab. 3 and occupied volume estimation in Tab. 3. Entirely dropping the ground-truth depth maps and only using the estimated depths proves to be the most successful approach and is the one we use in practice. We hypothesize that this is due to the slight smoothing present in the depth maps produced by Depth Anything v2 [31], which may prevent the model from overfitting to some specific shape features in perfect depth maps of some training samples.
Even though it is trained on synthetic data, this observation further confirms the generalizability of 3DC to real data, since in a practical scenario such as the real data of Tab. 3, ground-truth depth maps are not available.
@@ 4.5 Limitations and future work
In Fig. 7, we show failure cases of our method where the shapes of the objects are too complex for our volume occupancy estimator. Future work will focus on refining this estimate to handle highly non-convex and complex shapes. Additionally, and similarly to previous works [12, 15], our method currently expects the user to select the area of interest in at least one of the images. Automating this step as perfomed in [30] would enhance the ease of use of our method.
@@
Unlike our method, previous works additionally attempt to localize the objects being counted, increasing the interpretability of the results. However, these localizations are often erroneous when objects are stacked together, as illustrated in Fig. 9, greatly limiting their applicability. Thus a possible direction for further enhancements lies in integrating a robust localization of visible instances.
@@ 5 Conclusion
In this paper, we introduced a novel 3D counting method designed to address the challenges posed by stacks of identical objects, where occlusions and irregular arrangements make accurate counting difficult. By decomposing the counting task into two complementary subproblems—3D geometry estimation and occupancy ratio analysis—we demonstrated that our approach can achieve reliable counts in a wide variety of cases. Through extensive experiments on both real-world and large-scale synthetic datasets, we validated the robustness and accuracy of our method across diverse scenarios, highlighting its potential applications in domains such as logistics, monitoring, and scientific research.
Our pipeline combines geometric reconstruction with deep learning depth analysis, establishing a framework that can generalize well to objects in stacks, packaging, or in a container. By releasing our datasets to the public, we aim to encourage future research and development in 3D counting. In summary, our method represents a significant step forward in visual counting, especially in challenging, stacked 3D settings, and we believe it will inspire new approaches to handling occlusion and spatial complexity in vision counting tasks.
@@ 6 Dataset Details
We capture 20 real scenes where the objects to count can be any stack of items that are at least partially visible. This includes stacked objects on a table or on the floor, objects in containers such as bowls or boxes, or objects still in their packaging. In Fig. 10, we display all the scenes that were not already presented in our main submission.
@@
Initially, we experimented with triangulation methods using two pairs of corresponding points across images. This would enable the calculation of a 3D distance and allow us to scale the scene. However, this approach proved unstable, as small inaccuracies in point matching led to significant variations in the scaling. Instead, we reconstructed the 3D scene with 3DGS [7] and measured the 3D distance directly within the reconstruction. This measurement allows us to rescale the scene to match the reference measurement. Note that the 3D point cloud generated by COLMAP [21], which is used as an initialization by 3DGS, is also scaled accordingly.
@@
For each scene, we require the unit volume of the object being counted. For many common food items, such as kidney beans or corn, this information is readily available online. For other scenes, the volume can be approximated, for example in the case of the beads in Fig. 1, by subtracting the volume of a cylinder from that of a sphere. Alternatively, we determine the volume by measuring the number of units needed to displace 100 milliliters of water.
@@
Using this method, we capture 20 scenes consisting of various items in different environments. The scenes vary in complexity, from simple quasi-spherical objects in containers to more challenging configurations, such as complicated shapes still in packaging (e.g., in the pasta scene). Occasionally, we reuse items for a second scene in a different setting, such as by removing objects from their packaging or changing containers.
beans: Est. 161, G.T. 207
chickpeas: Est. 2668, G.T. 3276
chocolates: Est. 201, G.T. 131
crosses: Est. 40, G.T. 205
crosses2: Est. 317, G.T. 205
gums2: Est. 500, G.T. 397
popcorn: Est. 2229, G.T. 2921
puzzle: Est. 574, G.T. 1000
screws: Est. 295, G.T. 514
screws2: Est. 627, G.T. 494
semicrosses: Est. 156, G.T. 257
semicrosses2: Est. 196, G.T. 257
Figure 10: Additional dataset samples. We report the results on all real scenes that are not already present in our main submission.

### v2

@@ 4 Experiments
BMNet+ [21] 0.91 0.87 320.50 158.87
SAM+CLIP [7, 16] 0.73 0.61 259.22 102.77
CNN 0.66 0.48 235.74 98.44
ViT+H 0.42 0.24 149.90 47.36
Ours 0.22 0.09 79.48 27.65
Table 1: Counting evaluation on 100 synthetic scenes.
@@
BMNet+ [21] 0.93 0.98 966.76 131.44
SAM+CLIP [7, 16] 0.94 0.99 980.33 124.31
CNN 0.95 0.93 992.06 97.09
ViT+H 0.94 0.93 979.29 91.45
Human 0.79 0.84 823.23 76.85
Human-Vote 0.60 0.30 621.46 57.91
LlamaV 3.2 1.00 1.00 1037.5 190.48
Ours (Color) 0.57 0.27 607.98 74.33
Ours 0.36 0.06 382.59 53.31
Table 2: Counting evaluation on real-world scenes.
[…]
@@
Table 3: Occupancy ratio estimation. We evaluate our method against three additional baselines that are tasked with prediction the occupancy ratio $\gamma$ from a depth map.
@@ 4.1 Metrics
We use several metrics to assess the accuracy of object counting and occupancy ratio estimation. The object counts vary significantly across scenes, ranging from 19 to 20063. Thus, the Mean Absolute Error (MAE)—defined as $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|,$ where $y_{i}$ is the ground truth count and $\hat{y}_{i}$ is the predicted count for each scene—tends to amplify the importance of scenes with a high counts. To mitigate this, we also report normalized metrics. We use the Normalized Absolute Error (NAE), Squared Relative Error (SRE), and Symmetric Mean Absolute Percentage Error (sMAPE), which scale errors relative to the ground truth. The NAE provides a measure of the absolute error normalized by the total ground truth count across scenes, SRE emphasizes larger errors and penalizes significant deviations in high-count scenes, and , sMAPE offers a normalized percentage error. The exact formulas for all metrics can be found in our supplementary.
4.2 Counting Evaluation
(a) $\mathcal{N}_{est}=38$, $\mathcal{N}_{gt}=36$
(b) $\mathcal{N}_{est}=2133$, $\mathcal{N}_{gt}=1830$
(c) $\mathcal{N}_{est}=338$, $\mathcal{N}_{gt}=397$
(d) $\mathcal{N}_{est}=261$, $\mathcal{N}_{gt}=300$
Figure 6: Additional qualitative results.
(a) (a)
(b) (b)
(c) (c)
(d) (d)
(e) (a)
(f) (b)
(g) (c)
(h) (d)
(i) (a)
(j) (b)
(k) (c)
(l) (d)
Figure 8: Intermediate results. From images (a), we find the key viewpoint and compute its depth (b) to estimate the occupancy ratio $\gamma$. Using the unit shape previously reconstructed from images of a template (c), and the overall reconstruction (d), we deduce the final count. (Pasta: $\gamma_{est}=30.5\%$, $N_{est}=509$, $N_{gt}=588$ Bricks: $\gamma_{est}=31.8\%$, $N_{est}=73$, $N_{gt}=100$ Crosses: $\gamma_{est}=29.6\%$, $N_{est}=88$, $N_{gt}=116$ )
@@ 4.2 Counting Evaluation
As far as we know, there is no previous work on counting from multiple images which does not either implicitly assume all objects to be visible or require additional sensors such as LiDARs. Thus, we compare 3DC against BMNet+ [21]. It predicts a density map over all pixels of an image, and the estimated count is then inferred by summing over all pixels. Additionally, we compare against a combination of SAM [7] and CLIP [16], where SAM is used to generate a large number of masks from an input image and CLIP uses a set of negative and positive text prompts to identify masks that represent an object of interest. The final count is then taken to be the number of these masks. In Tables 3 and 3, we compare the performance of these two baselines against that of our method. We outperform them in both cases.
In early experiments, we attempted to directly predict object count from images. To this end, we trained different networks, coined ViT+H and CNN. We also report the results in Tables 3 and 3. They perform poorly, especially on the real-world dataset, which is what prompted us to look into decomposing the problem into occupancy and volume estimation.
We also sought to estimate how good humans are at this counting task. To this end, we organized a contest and encouraged participants to make accurate guesses on the 45 real scenes. The contest registered 1485 guesses from 33 participants. We define the Human baseline as the average of error metrics of participants, and Human-Vote as the error of the average guess across all participants. This second baseline should be stronger as the errors of participants tend to cancel-out and that is what we observe in Tab. 3. However, the results are still much worse than what our approach delivers. Interestingly, participants who spent a longer time did not perform better than their peers, highlighting the difficulty of this task. When asked about the method they used, a majority of participants reported counting the number of objects on each axis, and multiplying these values together. This proved to be ineffective, however.
Finally, we tested LlamaVision 3.2 11B, a Large Language and Vision Model, on our counting task. While it was very good at describing the physical appearance of the objects in the real scenes and the composition of the scene, its count estimates were completely off as shown in Tab. 3.
Our approach beats all the aforementioned baselines, providing the first method to estimate stack counts with reasonable accuracy. However, we also note that our perfomance remains better on the synthetic data than on our benchmark of real scenes. This can be attributed to the additional complexity of these real scenes, which often contain thousands of objects and can be more challenging for volume estimation.
4.3 Occupancy Ratio Evaluation
We now focus on the occupancy ratio network alone. There is little work on occupancy ratio estimation, we thus implemented additional baselines to gain insights about the inner working of our approach.
@@ 4.3 Occupancy Ratio Evaluation
Since we assume that the depth map contains enough information to predict the occupancy ratio, we define a first baseline that we will refer to as DepthExtrapolated. Given the top view of the container, we compute the maximal depth using a monocular depth estimator and use it to normalize the depth map. Then, we average the resulting values of the $K$ pixels, yielding the volume fraction estimate
$\gamma_{est}^{norm}=\frac{1}{K}\sum\frac{d_{i}}{d_{max}}\;.$
@@
This first baseline tends to predict values lower than expected. Thus, we defined a second one we dubbed DepthCorrected. It uses linear regression to correct $\gamma_{est}^{norm}$ into a new estimate $\gamma_{est}^{corrected}$. This yields a method able to model the observation that depth maps with high variance tend to correspond to low occupancy ratios. Finally, we also compare with the mean estimator, that predicts the mean percentage of $32.3\%$ occupancy ratio for all inputs.
@@
As can be seen in Tab. 3, our method outperforms these three baselines by a significant margin. DepthCorrected is better than the other two baselines, which shows that depth information is indeed useful for this task. However, it does not fully predict the occupancy ratio. We interpret this as a strong clue that the ratio $\frac{d_{i}}{d_{max}}$ alone is not enough to predict $\gamma$, and our proposed network successfully learns to extract meaningful information from the depth maps. We hypothesize that our volume network captures additional geometric information such as the influence of concavities in the final occupation of volume.
@@ 4.4 Ablation Study
NAE $\downarrow$ SRE $\downarrow$ MAE $\downarrow$ sMAPE $\downarrow$
($\mathcal{T}-$, $\mathcal{D}-$) 0.22 0.09 79.48 27.65
($\mathcal{T}+$, $\mathcal{D}-$) 0.28 0.11 100.12 30.93
($\mathcal{T}+$, $\mathcal{D}+$) 0.31 0.12 111.04 35.92
Table 4: Ablation study on 3D counting. If ground-truth depth maps are used during training, it is indicated as $\mathcal{T}+$, and $\mathcal{T}-$ otherwise. Similarly, for evaluation purposes if ground-truth depth-maps are used during validation, we indicate it as $\mathcal{D}+$.
MAE $\downarrow$ RMSE $\downarrow$ sMAPE $\downarrow$ $R^{2}$uparrow
($\mathcal{T}-$, $\mathcal{D}-$) 0.06 0.07 29.18 0.79
($\mathcal{T}+$, $\mathcal{D}-$) 0.08 0.11 32.01 0.52
($\mathcal{T}+$, $\mathcal{D}+$) 0.10 0.13 37.35 0.32
Table 5: Ablation study on occupancy ratio estimation.
We ran additional experiments to evaluate the sensitivity of our approach to the depth maps produced by the monocular depth estimator.
Since our synthetic dataset has ground-truth depth maps both for the training images and the validation images, we use them as follows. Recall from Sec. 3.2 that, at training time, we use Depth Anything V2 [30] depth maps, a setting we refer to as $\mathcal{T}-$ for training without ground-truth. Instead, we could use ground-truth depth maps during training, a setting we will refer to as $\mathcal{T}+$. Similarly, at inference time, we can use the estimated depth map. which is what we normally do, or the ground-truth one. We refer these as $\mathcal{D}-$ and $\mathcal{D}+$, respectively. Thus, the standard configuration of our method is $\mathcal{T}-,\mathcal{D}-$. The others are only used for ablation purposes.
We report the results in Tab. 5 and occupancy ratio estimation in Tab. 5. Entirely dropping the ground-truth depth maps and relying only on estimated ones, which is our standard operating procedure, is best. We hypothesize that this is due to the slight smoothing in the produced depth maps, which may prevent the model from overfitting to specific shape features in perfect depth maps.
Even though the network is trained on synthetic data, this observation further confirms the generalizability of 3DC to real data, since in a practical scenario such as the real data of Tab. 3, ground-truth depth maps are not available.
Finally, we also report in Tab. 3 an ablated method Ours (Color) where the $\gamma$-network takes an RGB image as input instead of a depth map. While this method still outperforms humans, it suffers from a significant performance drop and justifies the use of depth maps in our final approach.
@@ 4.5 Limitations and Future Work
Figure 9: Complexity Analysis. Each point represents the $\gamma$ occupancy ratio error for a shape in the validation set.
To evaluate the robustness of our method to complex shapes, we visualize in Fig. 9 the error in occupancy ratio estimation as a function of shape complexity. The latter is measured by summing a curvature complexity term with the ratio of the shape’s volume by the volume of its convex hull:
$C=\frac{\kappa}{\|x_{max}-x_{min}\|_{2}\kappa_{0}}+\frac{V_{hull}-V}{V_{hull}}$
where $\kappa$ is the integrated mean curvature of the shape, $\|x_{max}-x_{min}\|_{2}$ is a scaling factor and $\kappa_{0}$ is the maximum scaled $\kappa$ observed in the dataset. We only observe a slight error increase as the shapes become more complex,
Unlike ours, many earlier methods attempt to localize the objects being counted, increasing the interpretability and usability of the results. However, these localizations are often erroneous when objects are stacked together, as illustrated in Fig. 10, greatly limiting their applicability. Thus another possible direction for further enhancements lies in integrating a robust localization of visible instances, and an estimation of a possible configuration of invisible instances.
@@ 5 Conclusion
In this paper, we introduced a novel method to count sets of stacked nearly-identical objects. In this scenario, occlusions and irregular arrangements make accurate counting difficult. By decomposing the counting task into complementary subproblems—estimating independently the 3D volume of the stacks and the proportion of this volume actually occupied by objects—we were able to propose an effective solution that is easy to implement and far outperforms humans at this highly non-trivial task.
Our experiments show that the performance of our approach can degrade with increased geometric complexity or visually complex scenes. In future work, we will therefore look into training the volume occupancy estimator to overcome these challenges. More generally, we believe our method and proposed datasets will open new applications and encourage future works centered on stacks of 3D objects, including 3D reconstruction, counting, or 3D scene understanding.
@@ 6 Dataset Details
We capture 45 real scenes where the objects to count can be any stack of items that are at least partially visible. This includes stacked objects on a table or on the floor, objects in containers such as bowls or boxes, or objects still in their packaging.
@@
Initially, we experimented with triangulation methods using two pairs of corresponding points across images. This would enable the calculation of a 3D distance and allow us to scale the scene. However, this approach proved to be unstable, as small inaccuracies in point matching led to significant variations in the scaling. Instead, we reconstructed the 3D scene with 3DGS [6] and measured the 3D distance directly within the reconstruction. This measurement allows us to rescale the scene to match the reference measurement. Note that the 3D point cloud generated by COLMAP [20], which is used as an initialization by 3DGS, is also scaled accordingly.
@@
For each scene, we require the unit volume of the object being counted. For complex shapes, we determine the unit volume $v$ by 3D reconstruction, similar to the method described in our main paper. For many common food items, such as kidney beans or corn, this information is readily available online. For other scenes, the volume can be approximated, for example in the case of the beads in Fig. 1, by subtracting the volume of a cylinder from that of a sphere.
@@
Using this method, we capture 45 scenes consisting of various items in different environments. The scenes vary in complexity, from simple quasi-spherical objects in containers to more challenging configurations, such as complicated shapes still in packaging (e.g., in the pasta scene). Some items are used in several scenes, in which case the container and location are modified to create a new setting.
@@
7 Evaluation metrics
In this section, we provide the exact formula behind the metrics used in Sec. 4. As explained, the NAE and SRE are defined as:
NAE $=\frac{\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|}{\sum_{i=1}^{n}y_{i}},\hskip 14.40004pt\text{SRE}$ $=\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}y_{i}^{2}}\;.$
We also report the Symmetric Mean Absolute Percentage Error (sMAPE), which can be considered a normalized percentage error. It is expressed as follows:
The formula of sMAPE ensures that errors are scaled symmetrically between the prediction and ground truth counts. Finally, the coefficient of determination, $R^{2}$, measures the proportion of variance in the ground-truth occupancy ratio $\gamma$ explained by our predictions
[…]